# A minimalistic and general weighted averaging method for inconsistent data

### v2

@@ III.1 Synthetic tests
Figure 2: Standard and Jeffreys’ weighted averages of different simulated data sets: data randomly sampled from a normal distribution (top), and with the addition of a random bias (middle) or an outlier (bottom).
@@
    The first set of values $x_{i}$ is obtained by a normal distribution with a mean value of $\mu=1$ and a standard deviation equal to $\sigma=0.1$. For each data point, the uncertainty $\sigma_{i}=\sigma$ is considered.
@@
    The third set is the same as the first, but with the addition of an outlier at $\mu+5\,\sigma$ that has an uncertainty of $\sigma_{\mathrm{out}}=\sigma/3$.
@@
Figure 3: Final likelihood distribution for the data set 3 together with the input values.
@@
For the three sets, the standard weighted average, with or without Birge ratio correction, is compared to the conservative and Jeffreys’ weighted averages. The results are presented in Tables 13 and in Fig. 2 (for the standard and Jeffreys’ weighted averages only) and Fig. 3. As we can see, for normal and inconsistent data without outliers, the mean value is well reproduced by all methods. Because of the pessimistic priors on $\sigma_{i}^{\prime}$, the Jeffreys’ and conservative final uncertainties are generally larger than the standard uncertainty. For consistent data (set 1), they are larger by a factor of about two. As expected, for the inconsistent data (set 2), the uncertainty associated with the standard weighted average is significantly smaller than the others, with the Jeffreys’ uncertainty being the largest, followed by the conservative and the Birge uncertainties. Jeffreys’ weighted average uncertainty is almost six time larger than the standard weighted average uncertainty and double that of the Birge-ratio-corrected uncertainty.
@@
When an outlier is present, Jeffreys’ weighted average values are quite different from the standard weighted average. The effect of the presence of an outlier (set 3) is clearly visible in Fig. 3, where the final likelihoods are plotted together with the data, and in the results presented in in Table 3. As we can see, the effect on the standard likelihood, obtained by the product of Gaussian distributions, is drastic, with a shift in the direction of the outlier. If the data uncertainties are regarded as lower bounds only, the effect is greatly mitigated, resulting in just an asymmetry of the tails for the Jeffreys’ and conservative priors, which have very similar final probability distributions. This behaviour looks quite similar to other methods that utilize the deviation of the data point from the calculated value $\hat{\mu}$ to estimate the possible missing uncertainty contributions [3, 9]. However, unlike these past works, here the distributions are derived from the initial assumptions on the uncertainties, for which we consider a priori that the provided values $\sigma_{i}$ are only lower bounds of the real uncertainty.
III.2 The Newtonian constant of gravity
@@
A significant example of an average between independent and possibly inconsistent measurements is the determination of the Newtonian constant of gravity, which, due to the difficulties associated with its measurement, has long been the fundamental constant with the highest relative uncertainty. Such difficulties are mainly due to the challenging experimental conditions, where very small forces must be isolated from a noisy environment [19, 20]. The official value is provided by CODATA with a standard inverse-variance weighted average, and the associated uncertainty is eventually multiplied by an expansion factor to maintain consistency between the final result and the considered measurements.
@@
Figure 4: Comparison between the official CODATA values [21, 22, 23, 24, 25, 26, 27, 28, 29] of the Newtonian constant and the values obtained by the Bayesian weighted average using Jeffreys’ prior. CODATA values obtained from single measurements are presented alone, as no weighted average could be performed. The small error bar of the CODATA values indicates the uncertainty calculated by the standard weighted average, and the large one indicates the recommended uncertainty. The horizontal dashed line corresponds to the latest CODATA value (2022 edition [30], equal to the 2018 edition value).
@@
Here, we apply the proposed averaging method to all data sets included in the different editions of the CODATA compilation [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The results are presented together with the official values in Fig. 4. For each reported CODATA value, the large error bar corresponds to the recommended value of the uncertainty, and the small one to the uncertainty calculated by the standard weighted average. As we can see, the standard weighted average is, for some years, several standard deviations away from the most recent CODATA value from 2018 (the horizontal dashed line in the figure, which is the same as the more recent CODATA 2022 value [30]), considered here as the reference value. Contrary to standard procedures, one can see that the values obtained by the Jeffreys’ weighted average are consistently in good agreement, being less than one standard deviation away from the most recent CODATA value, and are characterised by a more plausible uncertainty.
@@
Table 4: Weighted average values with corresponding uncertainties for the data set of the CODATA 1998 compilation for the Newtonian gravitational constant and corresponding recommended value.
@@
Dose model average [13] 6.6746 0.0035
@@
Figure 5: Final likelihood distribution (in log-scale) of the measurements of the Newtonian constant included in the CODATA 1998 compilation [24]. The CODATA 1998 recommended value is also reported (in gray), which differs from the standard weighted average for the considered measurements (in red).
The 1998 case is particularly difficult due to the inconsistency within the data set, arising from one very precise measurement [31] that differs significantly from the average of the other measurements. The value was later found to be affected by a large systematic error [32]. Details of the analysis of this specific case are presented in Table 4 and Fig. 5. The CODATA recommended value is obtained by the standard weighted average of all values, excluding the suspicious measurement. The corresponding uncertainty is obtained by applying an expansion factor of 37 to the standard weighted average uncertainty to reflect the presence of the outlier. More precisely, the final uncertainty has been chosen to ensure that the difference between the recommended value and the outlier is four times larger than the final uncertainty. Like in set 3 of the previous section, the Jeffreys’ weighted average is only slightly affected by the outlier in this challenging case. Looking at Table 4, one can note that the Jeffreys’ weighted average has a significantly smaller uncertainty than the method tested on the same data set by Dose [13]. There, a Bayesian model average was applied on different likelihood functions.
@@ III.3 Particle properties
Another field that has to deal with very different measurements to compile reference data is particle physics. The Particle Data Group (PDG) [33], who is providing the official reference values, implements a very well documented procedure, mainly based on the Birge ratio and data selection. The goal of this selection is to minimise possible correlations between considered data and to exclude evident outliers. The muon magnetic anomalous moment is an exception. No measurements are selected except the one considered as unique reference from the most recent experiment [34] (which is now in agreement with the Standard Model predictions [35]).
Figure 6: Top: Comparison between the values obtained by the Bayesian weighted average using Jeffreys’ prior and the official PDG values of several particle properties [33]. All values are normalised to the PDG uncertainties. The two dashed horizontal lines correspond to $\hat{\mu}-\hat{\mu}^{\mathrm{PDG}}=\pm\sigma_{\hat{\mu}}^{\mathrm{PDG}}$, i.e. to ±$\pm 1$ in the relative scale. Bottom: Ratio between the uncertainties obtained for the Jeffreys’ weighted average and the official PDG uncertainties.
In this section, we evaluate the Jeffreys’ weighted average of some particle properties, and compare the results with the PDG recommended values. Moreover, to test the robustness of the Jeffreys’ weighted average, we evaluate two sets:
@@
Figure 8: Final likelihood distribution (in linear scale) corresponding to the PDG measurement selection for the proton radius. The PDG official value is also reported, which is very close to the standard weighted average. Due to the bi-modality of the final probability distribution, using a weighted average is completely inappropriate.
An extreme example of asymmetry and multimodality in the final probability distribution is the case of the proton rms charge radius. For this quantity, very different results are obtained from the Jeffreys’ weighted average with respect to the recommended PDG value and the standard weighted average (so different that they are not reported in Fig. 6). As we can see from the plot of the probability distribution (Fig. 8), a pronounced bi-modal distribution appears when evaluating the entire ensemble of available data. In this situation, any weighted average, for which we always assume unimodality of the associated probability distribution, is not suitable. This situation is similar to the case of a data set composed of only two data points with different values but the same uncertainties. Regardless of the chosen method, a weighted average will propose the midpoint between the data points due to the symmetry of the problem. In such cases, as written above, the whole probability distribution should be taken into account.
@@ IV The associated code
Despite the well-known problems of the standard weighted average based on the inverse of the variance, its still widespread use is undoubtedly largely due to the simplicity of its formula, which can be easily employed by anyone. The method presented here, as well as others proposed in the literature, has the disadvantage of requiring numerical methods for determining the weighted average and its uncertainty due to the complexity of the associated analytical formulas. The need for improved averaging methods is, however, prevalent, but due to the implementation difficulties, such alternative methods are generally quickly abandoned, sometimes in favour of the simpler Birge ratio. To close this gap, we provide a numerical tool for the proposed weighted averaging method. More precisely, we propose the Python library bayesian_average, which can be easily installed in any Python environment using the pip command and is also freely available via the GitHub repository  [^1] https://github.com/martinit18/bayesian_average. In addition to providing the weighted average based on the Jeffreys’ and conservative method priors, the standard inverse-variance method and its modified version with the Birge ratio are included for comparison. A graphical tool is available to plot the final weighted averages and the associated final likelihood probability distributions together with the input data. Figures 3, 5, 7 and 8 are typical outputs from bayesian_average (with minimal changes to the label of the axes).
@@ V Discussion and conclusions
We present a robust method for averaging inconsistent data, originally proposed by Sivia [14]. We suggest to use such a method as an alternative to the standard weighted average based on the inverse of the variance. In addition, we propose a Python package to easily calculate such a weighted average.
@@
Compared with other similar methods previously proposed in the literature, the number of working hypotheses is kept to the minimal requirements. The weighted average based on Bayesian statistics is proposed to avoid formulating complex hypotheses on the nature and behaviour of the unknown component of the true uncertainties. For each data point, a Gaussian distribution is considered, but the provided uncertainty $\sigma_{i}$ is regarded as a lower bound of the true uncertainty value. Using Bayes’ theorem and assuming a non-informative Jeffreys’ prior for $\sigma_{i}^{\prime}$, a new probability distribution is obtained by marginalising over $\sigma_{i}^{\prime}\in[\sigma_{i},\sigma_{i}^{\mathrm{max}}]$. The new arbitrary parameters $\sigma_{i}^{\mathrm{max}}$ are eliminated by looking at the asymptotic solution of the resulting weighted average $\hat{\mu}$. Contrary to the final associated probability distribution that diverges, the limit value of $\hat{\mu}$ and the associated uncertainty $\sigma_{\hat{\mu}}$ are well defined.
@@
The proposed method based on the Jeffreys’ prior is applied to a series of cases that show its reliability and robustness. For this purpose, simulated data, CODATA values of the Newtonian constant, as well as a series of particle property values from PDG are considered. In particular, for the CODATA 1998 case, where an outlier was causing issues, the method has proven to be a very robust tool. In the case of particle properties, different scenarios are encountered. For the largest part of the cases, the Jeffreys’ weighted average reproduces very well the PDG recommended values, but with a slightly larger uncertainty. In some cases, however, we show that a weighted average procedure should be taken with caution and the entire probability distribution should be considered instead, which is in agreement with the PDG recommendations. This is particularly true for the case of the proton radius, which shows a pronounced multimodality in the corresponding final probability. With the proposed method, these difficult cases are easily identified by looking at the final probability distribution.

### v3

@@ III.1 Synthetic tests
Figure 2: Standard and Jeffreys’ weighted averages of different simulated data sets: data randomly sampled from a normal distribution (top) and with the addition of a random bias (middle) or an outlier (bottom).
@@
    The first set of $N=20$ values $x_{i}$ is obtained by a normal distribution with a mean value of $\mu=1$ and a standard deviation equal to $\sigma=0.1$. For each data point, the uncertainty $\sigma_{i}=\sigma$ is considered.
@@
    The third set is the same as the first but with the addition of an outlier at $\mu+5\,\sigma$ that has an uncertainty of $\sigma_{\mathrm{out}}=\sigma/3$ with a total number of $N=21$ values.
The choice of $\sigma_{\mathrm{bias}}=10\,\sigma$ for set 2 reflects some typical measurement scenarios, such as those reported in Ref. Duval2024. For set 3, the $5\sigma$ separation is motivated by the five-sigma threshold usually implemented in particle physics. The value of $\sigma_{\mathrm{out}}=\sigma/3$ has been chosen arbitrarily to illustrate a clear deviation, but is still less extreme than the measurement discussed in Sec. III.2. The values of the sets are provided in the Supplementary data.
@@
Figure 3: Final likelihood distributions for the data set 3 together with the input values (black). The solid lines represent the standard (red), conservative (green) and Jeffreys’ prior likelihood distribution (blue). The corresponding average and standard deviation are marked by vertical lines.
@@
For the three sets, the standard weighted average, with or without Birge ratio correction, is compared to the conservative and Jeffreys’ weighted averages. The results are presented in Tables 13 and in Fig. 2 (for the standard and Jeffreys’ weighted averages only) and Fig. 3. As we can see, for normal and inconsistent data without outliers, the mean value is well reproduced by all methods. Because of the pessimistic priors on $\sigma_{i}^{\prime}$, the Jeffreys’ and conservative final uncertainties are generally larger than the standard uncertainty. For consistent data (set 1), they are larger by a factor of about two. As expected, for the inconsistent data (set 2), the uncertainty associated with the standard weighted average is significantly smaller than the others, with the Jeffreys’ uncertainty being the largest, followed by the conservative and the Birge uncertainties. Jeffreys’ weighted average uncertainty is almost six times larger than the standard weighted average uncertainty and double that of the Birge-ratio-corrected uncertainty.
@@
When an outlier is present, Jeffreys’ weighted average values are quite different from the standard weighted average. The effect of the presence of an outlier (set 3) is clearly visible in Fig. 3, where the final likelihoods are plotted together with the data, and in the results presented in Table 3. As we can see, the effect on the standard likelihood, obtained by the product of Gaussian distributions, is drastic, with a shift in the direction of the outlier. If the data uncertainties are regarded as lower bounds only, the effect is greatly mitigated, resulting in just an asymmetry of the tails for the Jeffreys’ and conservative priors, which have very similar final probability distributions. This behaviour looks quite similar to other methods that utilize the deviation of the data point from the calculated value $\hat{\mu}$ to estimate the possibility of missing uncertainty contributions Lira2007; Huang2018. However, unlike these past works, here the distributions are derived from the initial assumptions on the uncertainties, for which we consider a priori that the provided values $\sigma_{i}$ are only lower bounds of the real uncertainty.
III.2 The Newtonian constant of gravitation
@@
A significant example of an average between independent and possibly inconsistent measurements is the determination of the Newtonian constant of gravitation, which, due to the difficulties associated with its measurement, has long been the fundamental constant with the highest relative uncertainty. Such difficulties are mainly due to the challenging experimental conditions, where very small forces must be isolated from a noisy environment Rothleitner2017; Merkatas2019. The official value is provided by CODATA with a standard inverse-variance weighted average, and the associated uncertainty is eventually multiplied by an expansion factor to maintain consistency between the final result and the considered measurements.
@@
Figure 4: Comparison between the official CODATA values CODATA1969; CODATA1973; CODATA1986; CODATA1998; CODATA2002; CODATA2006; CODATA2010; CODATA2014; CODATA2018 of the Newtonian constant and the values obtained by the Bayesian weighted average using Jeffreys’ prior. CODATA values obtained from single measurements are presented alone, as no weighted average could be performed. The small error bar of the CODATA values indicates the uncertainty calculated by the standard weighted average, and the large one indicates the recommended uncertainty. The horizontal dashed line corresponds to the latest CODATA value (2022 edition CODATA2022, equal to the 2018 edition value).
@@
Here, we apply the Jeffreys’ weighted average to all data sets included in the different editions of the CODATA compilation CODATA1969; CODATA1973; CODATA1986; CODATA1998; CODATA2002; CODATA2006; CODATA2010; CODATA2014; CODATA2018; CODATA2022. The results are presented together with the official values in Fig. 4. For each reported CODATA value, the large error bar corresponds to the recommended value of the uncertainty, and the small one to the uncertainty calculated by the standard weighted average. As we can see, the standard weighted average is, for some years, several standard deviations away from the most recent CODATA value from 2018 (the horizontal dashed line in the figure, which is the same as the more recent CODATA 2022 value CODATA2022), considered here as the reference value. Contrary to standard procedures, one can see that the values obtained by the Jeffreys’ weighted average are consistently in good agreement, being less than one standard deviation away from the most recent CODATA value, and are characterised by a more plausible uncertainty.
@@
Table 4: Weighted average values with corresponding uncertainties for the data set of the CODATA 1998 compilation for the Newtonian gravitational constant and corresponding recommended value. All reported values are expressed in units of $10^{-11}$ m3 kg-1 s-2.
@@
Dose model average Dose2007 6.6746 0.0035
@@
Figure 5: Final likelihood distribution (in log scale) of the measurements of the Newtonian constant included in the CODATA 1998 compilation CODATA1998. The CODATA 1998 recommended value is also reported (in grey), which differs from the standard weighted average for the considered measurements (in red).
The 1998 case is particularly difficult due to the inconsistency within the data set, arising from one very precise measurement Michaelis1995 that differs significantly from the average of the other measurements. The value was later found to be affected by a large systematic error Michaelis2004. Details of the analysis of this specific case are presented in Table 4 and Fig. 5. The CODATA recommended value is obtained by the standard weighted average of all values, excluding the suspicious measurement. The corresponding uncertainty is obtained by applying an expansion factor of 37 to the standard weighted average uncertainty to reflect the presence of the outlier. More precisely, the final uncertainty has been chosen to ensure that the difference between the recommended value and the outlier is four times larger than the final uncertainty. Like in set 3 of the previous section, the Jeffreys’ weighted average is only slightly affected by the outlier in this challenging case. Looking at Table 4, one can note that the Jeffreys’ weighted average has a significantly smaller uncertainty than the method tested on the same data set by Dose Dose2007. There, a Bayesian model average was applied on different likelihood functions.
@@ III.3 Particle properties
Another field that has to deal with very different measurements to compile reference data is particle physics. The Particle Data Group (PDG) PDG2024, which is providing the official reference values, implements a very well-documented procedure, mainly based on the Birge ratio and data selection. The goal of this selection is to minimise possible correlations between considered data and to exclude evident outliers. The muon magnetic anomalous moment is an exception. No measurements are selected except the one considered as unique reference from the most recent experiment Aguillard2023 (which is now in agreement with the Standard Model predictions Ignatov2024).
Figure 6: Top: Comparison between the values obtained by the Bayesian weighted average using Jeffreys’ prior and the official PDG values of several particle properties PDG2024. All values are normalised to the PDG uncertainties. The two dashed horizontal lines correspond to $\hat{\mu}-\hat{\mu}^{\mathrm{PDG}}=\pm\sigma_{\hat{\mu}}^{\mathrm{PDG}}$, i.e. to $\pm 1$ in the relative scale. Bottom: Ratio between the uncertainties obtained for the Jeffreys’ weighted average and the official PDG uncertainties.
In this section, we evaluate the Jeffreys’ weighted average of some particle properties and compare the results with the PDG recommended values. Moreover, to test the robustness of the Jeffreys’ weighted average, we evaluate two sets:
@@
Figure 8: Final likelihood distribution (in linear scale) corresponding to the PDG measurement selection for the proton radius. The PDG official value is also reported, which is very close to the standard weighted average.
An extreme example of asymmetry and multimodality in the final probability distribution is the case of the proton rms charge radius. For this quantity, very different results are obtained from the Jeffreys’ weighted average with respect to the recommended PDG value and the standard weighted average (so different that they are not reported in Fig. 6). As we can see from the plot of the probability distribution (Fig. 8), a pronounced bi-modal distribution appears when evaluating the entire ensemble of available data. In this situation, any weighted average, for which we always assume unimodality of the associated probability distribution, is not suitable. This situation is similar to the case of a data set composed of only two data points with different values but the same uncertainties. Regardless of the chosen method, a weighted average will propose the midpoint between the data points due to the symmetry of the problem. In such cases, as written above, the whole probability distribution should be taken into account as well as the critical judgement of experienced researchers.
@@ IV The associated code
Despite the well-known problems of the standard weighted average based on the inverse of the variance, its widespread use is undoubtedly largely due to the simplicity of its formula, which can be easily employed by anyone. The Jeffreys’ weighted average, as well as other alternative methods proposed in the literature, has the disadvantage of requiring numerical methods for determining the weighted average and its uncertainty due to the complexity of the associated analytical formulas. The need for improved averaging methods is, however, prevalent, but due to the implementation difficulties, such alternative methods are generally quickly abandoned, sometimes in favour of the simpler Birge ratio. To close this gap, we provide a numerical tool for the proposed weighted averaging method. More precisely, we propose the Python library bayesian_average, which can be easily installed in any Python environment using the pip command and is also freely available via the GitHub repository  [^1] https://github.com/martinit18/bayesian_average. In addition to providing the weighted average based on the Jeffreys’ and conservative method priors, the standard inverse-variance method and its modified version with the Birge ratio are included for comparison. A graphical tool is available to plot the final weighted averages and the associated final likelihood probability distributions together with the input data. Figures 3, 5, 7 and 8 are typical outputs from bayesian_average (with minimal changes to the label of the axes).
@@ V Discussion and conclusions
We present a robust method for averaging inconsistent data, originally proposed by Sivia Sivia1996b. We suggest using this method for the analysis of simple data sets exhibiting inconsistencies as an alternative to the standard weighted average based on the inverse of the variance. In addition, we propose a Python package to easily calculate such a weighted average.
@@
Compared with other similar methods previously discussed in the literature, the number of working hypotheses is kept to the minimal requirements. The weighted average formulated by Sivia Sivia1996b and based on Bayesian statistics is implemented to avoid formulating complex hypotheses on the nature and behaviour of the unknown component of the true uncertainties. For each data point, a Gaussian distribution is considered, but the provided uncertainty $\sigma_{i}$ is regarded as a lower bound of the true uncertainty value. Using Bayes’ theorem and assuming a non-informative Jeffreys’ prior for $\sigma_{i}^{\prime}$, a new probability distribution is obtained by marginalising over $\sigma_{i}^{\prime}\in[\sigma_{i},\sigma_{i}^{\mathrm{max}}]$. The new arbitrary parameters $\sigma_{i}^{\mathrm{max}}$ are eliminated by looking at the asymptotic solution of the resulting weighted average $\hat{\mu}$.
Because of the independent treatment for each datum uncertainty $\sigma_{i}$, the approaches presented in the previous sections are well adapted to interlaboratory studies. Each uncertainty $\sigma_{i}$ may be underestimated due to unknown systematic errors specific to the corresponding laboratory.
A similar approach has been implemented by Mana Mana2021, using a prior on $\sigma^{\prime}_{i}$ as in Eq. (8). Unlike our proposed approach, Mana assumes that for some cases $\sigma_{i}=\sigma^{\prime}_{i}$ holds. To determine which data satisfy this assumption, the probability of the hypothesis (i. e., that $\sigma_{i}=\sigma^{\prime}_{i}$ is valid) is evaluated by Bayesian model selection, based on the marginal likelihood (Bayesian evidence). Related approaches are discussed in the literature Toman2012; Rukhin2019, where an unknown bias is considered for clusters of data, thereby leading to an additional uncertainty being added to the original one. Other similar methods are proposed in Refs. Shirono2010; Huang2018.
All of the above-mentioned implementations imply, however, introduce significant complexity and may reduce the generality of their application. In contrast, although more pessimistic (implicitly treating all laboratories with caution), our proposed method is much simpler and more generally applicable, while remaining well suited for data obtained from the same measurement apparatus.
The method based on the Jeffreys’ prior is applied to a series of cases that show its reliability and robustness. For this purpose, simulated data, CODATA values of the Newtonian constant, as well as a series of particle property values from PDG are considered.
In particular, in the presence of outliers, the method has proven to be a very robust tool. In the case of the synthetic data set 3 and the average of the Newtonian gravitational constant reported in CODATA 1998, the effect of the outlier is minimized and the reference value is recovered without problems. Nonetheless, the treatment of data containing outliers is always delicate. As an example, in the case of the proton radius, the 2010 measurement using muonic hydrogen Pohl2010 was considered an outlier and thus excluded from the CODATA compilation for several years CODATA2010; CODATA2014. Over time, however, this “anomalous” experiment motivated new measurements that highlighted systematic effects in past measurements not based on muonic hydrogen, leading to its inclusion in the CODATA compilation CODATA2022. For such cases involving outliers, as best practice, we suggest clearly detailing the data analysis choices and justifications of the chosen analysis method. This transparency ensures that these decisions are understandable and allows for the possible application of alternative approaches as new information or methods arise. For such cases, the standard weighted average method should, however, be always avoided, as it does not account for potential misestimation of uncertainties.
@@
In the case of particle properties, different scenarios are encountered. For the largest part of the cases, the Jeffreys’ weighted average reproduces very well the PDG recommended values, but with a slightly larger uncertainty. In some cases, however, we show that a weighted average procedure should be taken with caution, and the entire probability distribution should be considered instead, which is in agreement with the PDG recommendations. This is particularly true for the analysis of the latest compilation of the proton radius, which shows a pronounced multimodality in the corresponding final probability. With the presented method, these difficult cases are easily identified by looking at the final probability distribution.
Once more, it is important to emphasise that the Jeffreys’ weighted average should not be regarded as a substitute for more thorough, expert analysis. The critical judgement of experienced researchers is irreplaceable. However, this method serves as a valuable tool for obtaining simple weighted averages from inconsistent data, especially in cases where the commonly used standard weighted average may lead to misleading results.
@@
Acknowledgements.
We would like to thank François Nez for the very useful discussions on CODATA values, Louis Duval for providing the motivation to start this work with his analysis of “inconsistent” data, Mark Plimmer for the careful reading of the manuscript, Andrew Charman for introducing us to Sivia’s earlier work that confirmed our independent calculations, and Udo von Toussaint for his valuable suggestions.